OpenStudy (bluebeta):

A 50-foot tree casts a shadow 75 feet long. How can you find sine for the angle of elevation that is determined by the sun and the top of the tree? My answer is below.

OpenStudy (bluebeta):

c^2=502+752 c^2=5625+2500 c^2=8125 c=90.14 sin(θ)=50/90.14 sin(θ)=0.55

OpenStudy (bluebeta):

Is that correct? @phi @radar

OpenStudy (bluebeta):

50^2+75^2**

OpenStudy (mathstudent55):



OpenStudy (bluebeta):

yes i had that.

OpenStudy (mathstudent55):

\(50^2 + 75^2 = c^2\) \(c = 90.14 ~ft\) You got that right.

OpenStudy (mathstudent55):



OpenStudy (bluebeta):

okay i'm following

OpenStudy (mathstudent55):

\(\sin \theta = \dfrac{opp}{hyp} = \dfrac{50}{90.14} = 0.55\) You got that right too. Good job!

OpenStudy (bluebeta):

Yay!! So I DID IT :D lol mini proud moment :P Thank you for checking!!

OpenStudy (mathstudent55):

You're welcome. You should be proud for getting a correct answer, and also, your approach to OS is excellent. You solved a problem and showed everything you did, and just needed to know if it was correct. Good going!

OpenStudy (bluebeta):

Thank you very much!!

OpenStudy (mathstudent55):

You're welcome.